Which of the following sentences is a proposition whose truth value is T?


Select one:

a. x + 2 = 11

b. x + y = y + x for every pair of real numbers x and y

c. Answer this question carefully.

d. Manila is in Cagayan.




1
Expert's answer
2022-02-06T17:00:08-0500

a)a)

x+2=11.x + 2 = 11.

The given statement can neither be true or false. It depends on the value of x. Hence, it is not a proposition.


b)b)

x + y = y + x for every pair of real numbers x and y

This is a proposition whose truth value is T.T.

c)c)

 Answer this question carefully.

Since " Answer this question carefully" is not a statement that can be either true or false, it is not a proposition.


d)d)

Manila is in Cagayan.

This statement is a proposition whose truth value is FF .

Therefore, the correct choice is b.